OpenStudy (anonymous):

find two positive numbers whose sum is 17 and whose product is a maximum.

OpenStudy (anonymous):

The numbers are x and 17-x The product gives us: f(x)=x(17-x) = -x^2+17x To find the maximum you have to take the derivative of the function (-x^2+17x) and then solve it for zero. d/dx=-2x+17 d/dx=0 gives us -2x+17=0 x=17/2 <==> x=8.5 If you need the point of the maximum you insert x into f(x)=x(17-x) and get: f(x)=72.25 Hope this helps! //Ali

OpenStudy (anonymous):

very much so thanks